If 5x - 17 = -x + 7, then x =
Answer by lenny460(1073)

If 5x - 17 = -x + 7, then x =
Add 1x to each side of the equation.
5x - 17 + 1x = -1x + 7 + 1x
Combine like terms:
5x + 1x = 6x
-1x + 1x = 0
We now have:
6x - 17 = 7
Add 17 to each side of the equation.
6x - 17 + 17 = 7 + 17
Combine like terms:
-17 + 17 = 0
7 + 17 = 24

Therefore:
6x = 24
Divide each side by 6
6x/6 = 24/6
x = 4

Answer:
x = 4






Lennox Obuong
Algebra Tutor
Nairobi, Kenya
Email: obuong3@aol.com